OpenStudy (anonymous):

Find the slope of the line. 3x+5y=2

OpenStudy (anonymous):

The standard form of line is given by : \(ax + by = c\), for this Slope is given by: \[Slope = - \frac{ x \; \text{coefficient}}{y \; \text{coefficient}}\]

OpenStudy (anonymous):

What is x-coefficient there can you tell?

OpenStudy (anonymous):

x-coefficient is the term that is attached to \(x\), what is it?

OpenStudy (anonymous):

Do you believe in replying or interacting?? :P

OpenStudy (anonymous):

sorry! is it 3

OpenStudy (anonymous):

And for y it is?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

So just plug in : \[Slope = -\frac{3}{5}\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Thats it? like thats the answerr? lol

OpenStudy (anonymous):

\(\text{This is it..}\)

OpenStudy (anonymous):

But point to remember is : you equation should be in Standard Form, this formula is applicable for Standard Form only.. :)

OpenStudy (anonymous):

*your

OpenStudy (anonymous):

oh okay i thought it was much harder then that lol thank you

OpenStudy (anonymous):

Standard Form looks like this : \(ax + bx = c\), like your equation : 3x + 5y = 2

OpenStudy (anonymous):

*\(by\)

OpenStudy (anonymous):

Did you get this?

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

Okay, good..!! :)

OpenStudy (anonymous):